OpenStudy (anonymous):

help please(: 1 over 3 (2x − 8) = 4

OpenStudy (anonymous):

\[\frac{1}{3(2x-8)}=4\] What do you suggest from here?

OpenStudy (hba):

\[\frac{ 1 }{ 3(2x-8) }=4\] Use distributive property to evaluate your denominator

OpenStudy (anonymous):

@Limitless its 1 over 3 and next to (2x − 8) = 4

OpenStudy (anonymous):

\[\frac{1}{3}(2x-8)=4\] What do you suggest from here?

OpenStudy (anonymous):

@Limitless distributing?

OpenStudy (hba):

\[\frac{ 1 }{ 3 }(2x-8)=4\] 1)Multiply both sides by 3

OpenStudy (anonymous):

@sylinan Sure. Then what?

OpenStudy (anonymous):

@hba doesnt it cancel out?

OpenStudy (hba):

Yes.

OpenStudy (anonymous):

You get \[\frac{2}{3}x-\frac{8}{3}=4.\]

OpenStudy (hba):

@Limitless Stop complicating it.

OpenStudy (hba):

Multiply your question by 3

OpenStudy (anonymous):

@hba Complicating it would be using the Taylor series for \(x\). I apologize for replying when she was talking to you; I misread the post.

OpenStudy (hba):

@sylinan \[\frac{ 1 }{ 3 }(2x-8)=4\] Multiply both sides by 3.

OpenStudy (hba):

\[3 \times\frac { 1 }{ 3 }(2x-8)=4 \times 3\]

OpenStudy (hba):

@sylinan I know you can do it now :)

OpenStudy (anonymous):

@hba okay so far i got 3(2x-8)=12

OpenStudy (hba):

Now see this \[\ \cancel 3 \times \frac{ 1 }{\ \cancel 3 }(2x-8)=4 \times 3\]

OpenStudy (anonymous):

so then its 1(2x-8)=12?

OpenStudy (hba):

Now add both sides by 8

OpenStudy (anonymous):

i thought i was suposed to do distribute the one?

OpenStudy (hba):

It doesn't make a diff It will stay the same :)

OpenStudy (anonymous):

true okay so i got x=10?

OpenStudy (hba):

Yeaahhhh :D

OpenStudy (hba):

You are right :P

OpenStudy (anonymous):

thank you (: help me with one more?(:

OpenStudy (hba):

Sure,Please post your question in a new tab :)

OpenStudy (anonymous):

okay so open a new question?

OpenStudy (hba):

Yeah. The CoC says that not me :P

OpenStudy (anonymous):

oh okay (: